Carol is thinking of a number. If the number is increased by 6, then doubled, and 9 is subtracted from the product, the result is 23. What is the original number?
Answer by ReadingBoosters(3246)

Visualize this problem as working backward like following a map in reverse from the destination back to the starting place.
[(23 + 9)/2]-6
32/2
16 - 6
10
Proof
10+6 = 16 *2 = 32 - 9 = 23